# A Simulation Based Inference Approach to the Dynamics of the MW–LMC System - Validation

### v1

@@ 1 Introduction
The Milky Way (MW) is undergoing a merger with the LMC [^1] The Large Magellanic Cloud. (see Vasiliev, 2023, for a comprehensive review detailing the effects of the LMC on the MW). The LMC is thought to be on its first pericentric passage and to have a dark matter mass $M_{\mathrm{LMC}}\sim 10^{11}\,\mathrm{M}_{\odot}$ (Besla et al., 2007, 2010; Boylan-Kolchin et al., 2011; Peñarrubia et al., 2016; Kravtsov & Winney, 2024). An alternative scenario has the LMC on its second pericentric passage, although most features of the earlier passage are superseded by the more recent passage at a smaller pericenter (Vasiliev, 2024). Many Local Group phenomena require a large mass for the LMC to explain: for example, the kinematics of its globular clusters (Watkins et al., 2024), the kinematics of MW satellites (Correa Magnus & Vasiliev, 2022; Kravtsov & Winney, 2024); dynamical models of MW stellar streams (Erkal et al., 2019; Koposov et al., 2019; Shipp et al., 2021; Vasiliev et al., 2021; Koposov et al., 2023); and the timing argument (Peñarrubia et al. 2016, but see also Benisty et al. 2022; Chamberlain et al. 2023; Benisty 2024) all require an LMC mass $M_{\mathrm{LMC}}\sim 10-20\,\times 10^{10}\,\mathrm{M}_{\odot}$ (see fig. 1 of Vasiliev, 2023, for a summary of LMC mass estimates). The orbit of the LMC is sensitive to the assumed Galactic potential (see fig. 3 of Vasiliev, 2023) and, because the LMC is of considerable mass, it is also subject to dynamical friction from the MW dark matter halo (Chandrasekhar, 1943).
@@
At present-day the LMC is at a Galactocentric distance $d=49.6\pm 0.5\,\mathrm{kpc}$ (Pietrzyński et al., 2019), and heliocentric line of sight velocity of $v_{\mathrm{los},}=262.2\pm 3.4\,\mathrm{km}\,\mathrm{s}^{-1}$ (van der Marel & Kallivayalil, 2014). There is more ambiguity in the measurement of the position of the LMC centre (see sec. 4.1, van der Marel & Kallivayalil, 2014) which correspond to different reported values of the mean proper motions (Gaia Collaboration et al., 2018, 2021; Wan et al., 2020) due to both perspective effects and internal motion in the LMC. Table. 2 in Vasiliev (2023) gives an overview of some of the recent measurements of the mean LMC proper motions and the associated centre. Although, more recently, hyper velocity stars that are ejected via interactions with the LMC’s supermassive black hole have been used to constrain both the LMC’s dynamical centre and past orbit (Han et al., 2025, Lucchini et al. in prep).
@@
As the LMC, a satellite galaxy, falls into the gravitational potential of the MW, the central galaxy, responds by generating a density wake (Chandrasekhar, 1943). This is because an infalling satellite will have a broad range of orbital frequencies that resonate with dark matter particles of the host galaxy (Mulder, 1983; Weinberg, 1986). The classical ‘conic’ wake trailing the LMC is described as the transient response, whereas the response elsewhere in the MW halo is the collective response caused by the barycentre displacement (Garavito-Camargo et al., 2019, 2021a; Tamfal et al., 2021; Foote et al., 2023). The density wake of the LMC is predicted to leave an observable signature in the density and kinematics of MW halo stars (e.g., Conroy et al., 2021; Cavieres et al., 2024; Chandra et al., 2025; Yaaqib et al., 2024; Byström et al., 2024; Amarante et al., 2024).
@@
As the LMC is massive, $M_{\mathrm{LMC}}\sim 10-20\,\times 10^{10}\,\mathrm{M}_{\odot}$ (a 1 : 5–10 mass ratio with the MW), and it has just passed its pericentre with a relative velocity $>300\,\mathrm{km}\,\mathrm{s}^{-1}$ (Hunt & Vasiliev, 2025), this has caused significant dynamical disequilibrium in the Galaxy. Specifically, the inner and outer parts of the MW halo have experienced different strengths of gravitational acceleration due to the LMC. To the Galactocentric observer, this appears as if the Galactic Northern sky is red-shifted and the Southern sky is blue-shifted as the halo moves preferentially ‘up’, towards the Galactic North. This displacement, known as the ‘reflex motion’, manifests itself as a dipole signal in density (Belokurov et al., 2019; Garavito-Camargo et al., 2021b; Conroy et al., 2021; Amarante et al., 2024) that is higher in the Galactic North, and similarly there exists a dipole in stellar radial velocities (Erkal et al., 2019; Petersen & Peñarrubia, 2020, 2021; Erkal et al., 2021; Yaaqib et al., 2024; Chandra et al., 2025; Byström et al., 2024). The magnitude of the velocity dipole is called the travel velocity, $v_{\rm{travel}}$, and its orientation is called the apex direction of the reflex motion $(l_{\rm{apex}}$, $b_{\rm{apex}})$ in Galactic coordinates. Although recent studies agree that the travel velocity direction and magnitude should point towards a point on the past orbit of the LMC, they do not converge on their values (e.g., Byström et al., 2024, fig. 9). This is likely due to each study using different stellar tracers over varying radial ranges and sky coverages.
@@
The MW–LMC system can be modelled at varying levels of simulation fidelity. There exists a trade off between the simulation fidelity and the ability to explore large model parameter spaces (e.g., see, Vasiliev et al., 2021, in the context of modelling the Sagittarius stream). The simplest prescription are rigid models of the MW and LMC galaxies. Rigid models describe the MW and LMC as analytic potentials that have fixed functional forms i.e., they are time-invariant, although they are allowed to move in response to each other. The main advantage of these models is their computational inexpense ($\sim\mathcal{O}(\rm seconds)$). However, the main drawback is their inability to truly capture the physics of galaxy formation e.g., the galaxy masses will change over time, which changes the gravitational potential, and hence their resulting interaction. To accommodate many further complexities, the MW–LMC system can instead be modelled using $N$-body simulations in combination with Basis Function Expansions (BFEs, e.g., Lilley et al., 2018a, b; Sanders et al., 2020; Petersen & Peñarrubia, 2020; Garavito-Camargo et al., 2019, 2021b; Lilleengen et al., 2023; Vasiliev, 2024). These models are often labelled as idealised or deforming MW–LMC simulations. They aim to match the present-day conditions of the MW and LMC while also accounting for the deformations to both galaxies due to the re-distribution of dark matter due to the LMC’s infall. This includes the formation of a dark matter dynamical friction wake trailing the LMC (Chandrasekhar, 1943). These deformations are not something that the simpler rigid MW–LMC models can capture although the dynamical friction can still be accounted for. However, they are significantly more computationally expense to run ($\sim\mathcal{O}(\rm months)$) as they contain billions of particles. As such, the parameter space explored for MW and LMC properties is much smaller e.g., only a handful of models exist to vary the LMC’s mass (Garavito-Camargo et al., 2019). Further, there have been MW–LMC pairs identified within hydrodynamical cosmological simulations. For example, there are many cases within the Feedback In Realistic Environments (FIRE, Wetzel et al., 2023) Latte, the Milky Way-est (Buch et al., 2024), the DREAMS (Rose et al., 2025) and the Auriga simulation suites (Grand et al., 2017, 2024). This is by far the most physically realistic setting to study MW–LMC interactions. However, in a similar vein to deforming simulations, these cosmological simulations can be computationally expensive and cover somewhat limited model parameter spaces.
@@
Simulation Based Inference (SBI, see, Cranmer et al., 2020, for a recent review) offers a route to strike a balance between simulation fidelity and the exploration of model parameter space. SBI is a powerful statistical framework for performing inference in complex modelling scenarios where traditional analytic methods are impractical or impossible. It is particularly well-suited for estimating parameter posterior distributions when the likelihood function cannot be explicitly defined, or is intractable from simulation. Instead, forward simulations are used to generate samples of the data and parameters. This technique is also known as ‘likelihood free inference’ or ‘distribution free inference’. Although, these names are slightly misleading as the inference is not likelihood-free, rather we avoid explicitly defining a likelihood and instead model it using many forward simulations. For our application to the dynamics of the MW–LMC interaction, removing the requirement to provide a well-defined likelihood is beneficial as we are no longer biased by assuming some functional form e.g., a Gaussian distribution for the MW reflex parameters: $v_{\rm{travel}}$, $l_{\rm{apex}}$, and $b_{\rm{apex}}$ (Petersen & Peñarrubia, 2021; Yaaqib et al., 2024; Byström et al., 2024; Chandra et al., 2025). This is especially important in the context of cosmological MW–LMC analogues as both the model parameter space and the posteriors can be non-trivial. SBI has been applied across many fields of research: from mathematics (e.g., Verdier et al., 2022), to quantitative finance (e.g., Yu, 2012), to seismology (e.g., Saoulis et al., 2024), therefore owing itself to become an effective lingua franca for facilitating communication across scientific domains. Furthermore, SBI has been increasingly used throughout astrophysics problems (e.g., Weyant et al., 2013; Alsing et al., 2019; Jeffrey et al., 2021; Lemos et al., 2021; Hermans et al., 2021; von Wietersheim-Kramsta et al., 2024; Lovell et al., 2024; Widmark & Johnston, 2025; Sante et al., 2025; Xiangyuan Ma et al., 2025; Jeffrey et al., 2025).
@@
For the first time, we will demonstrate the validity of using SBI to infer properties of the MW–LMC system when using measured dynamical quantities, i.e., the reflex motion, from many forward models the MW–LMC system using rigid prescriptions. We will investigate the ability of our SBI framework to infer the true parameter values of MW–LMC simulations described by rigid, deforming and cosmological models. This will demonstrate the ability that SBI has to upscale the fidelity of simpler simulations to more complex ones. Ultimately, this allows more a rapid exploration of large model parameters spaces at the fraction of the computational cost required to run the higher fidelity simulations. We will show that our SBI framework is reliable and unbiased through a series of diagnostic tests. Our novel approach will be applied to real data with the aim of constraining properties of the MW, LMC and the dynamics of the induced MW reflex motion.
@@ 2 Simulations
This section provides the details for our low fidelity rigid MW–LMC simulations. Throughout this work, conversions between the heliocentric and Galactocentric cartesian frames use the right-handed Cartesian coordinate system with the Sun positioned at $\vec{r}_{\odot}=(-8.3,0.0,0.02)\,\rm{kpc}$ (GRAVITY Collaboration et al., 2019; Bennett & Bovy, 2019), with velocity $\vec{v}_{\odot}=(11.1,244.24,7.24)\,\rm{km}\,\rm s^{-1}$ (Schönrich et al., 2010; Eilers et al., 2019).
@@ 2.1 The Milky Way – LMC potentials
We use the galaxy dynamics C++/Python package agama (Vasiliev, 2019) to generate many rigid MW–LMC simulations, each with a unique combination of model parameters. We describe the models for the MW and LMC in the following sections.
@@ 2.1.1 The Milky Way
Our MW model consists of a spherical bulge, an exponential stellar disc, and a dark matter halo. We keep the parameters of the stellar distributions fixed and close to the commonly used values suggested by McMillan (2017). The bulge density profile is:
@@
with a total mass $5\times 10^{10}\,\mathrm{M}_{\odot}$, a scale radius $R_{d}=3\,\mathrm{kpc}$, and a scale height $h_{d}=0.4\,\mathrm{kpc}$. We use a Navarro-Frenk-White (NFW, Navarro et al., 1996, 1997) dark matter halo density profile described by $M_{200,\rm MW}$ and $c_{200}$. These quantities are defined by a sphere enclosing an overdensity that is 200 times the critical density of the Universe, $\rho_{\mathrm{crit}}=3H^{2}/8\pi G$, as denoted by the ‘$200$’ subscript. We constrain the normalisation of the halo mass profile such that the circular velocity at the solar position is approximately $235\,\mathrm{km}\,\mathrm{s}^{-1}$ (e.g., matching constraints from McMillan, 2017, within the associated uncertainty). In combination with a chosen $M_{200,\rm MW}$ value, this allows us to find the corresponding concentration parameter, $c_{200}$.
@@ 2.1.2 The LMC
[…]
@@ 2.1.3 The MW–LMC interaction
where $\mathbf{x}_{i}$, $\mathbf{v}_{i}$ and $\Phi_{i}$ (where $i=\{\mathrm{MW},\mathrm{LMC}\}$) are the position vectors, velocity vectors and the gravitational potentials of the MW and LMC in an inertial frame, respectively. The extra acceleration term, $\mathbf{a}_{\mathrm{DF}}$, accounts for Chandrasekhar dynamical friction on the trajectory of the LMC (Chandrasekhar, 1943; Binney & Tremaine, 2008; Jethwa et al., 2016),
@@
where $X=v_{\rm LMC}/\sqrt{2}\sigma_{\rm MW}$ and $\sigma_{\rm MW}$ and $\rho_{\rm MW}$ are the velocity dispersion and total density field of the MW. Following Vasiliev et al. (2021), we take a fixed value of $\sigma_{\rm MW}=120\,\rm km\,\rm s^{-1}$ for the velocity dispersion as the dynamical friction is insensitive to the precise value. For the Coulomb logarithm we adopt $\ln\Lambda=\ln\left(100\,\rm kpc/\epsilon\right)$. The softening length, $\epsilon$, depends on the satellite’s density profile (White, 1976). We adopt $\epsilon=1.6\,a_{\rm LMC}$ as this has been used previously when modelling the LMC as a Plummer sphere (e.g., Hashimoto et al., 2003; Besla et al., 2007; van der Marel et al., 2012; Sohn et al., 2013; Kallivayalil et al., 2013). The numerator in the Coulomb logarithm expression an arbitrarily chosen value that loosely describes the average separation of the MW and LMC. In principle this value could be updated through the integration of the LMC orbit, however this will have a small effect. Furthermore, our use of a dimensionless parameter, $\lambda_{\mathrm{DF}}$, that modulates the strength of the dynamical friction that the LMC experiences will take into account changes to the fixed Coulomb logarithm value per simulation. Finally, we include the acceleration due to the reflex motion of the MW’s centre of mass towards the LMC.

### v2

@@ 1 Introduction
The Milky Way (MW) is undergoing a merger with the LMC [^1] The Large Magellanic Cloud. (see Vasiliev2023, for a comprehensive review detailing the effects of the LMC on the MW). The LMC is thought to be on its first pericentric passage and to have a dark matter mass $M_{\mathrm{LMC}}\sim 10^{11}\,\mathrm{M}_{\odot}$ (Besla2007; Besla2010; Boylan-Kolchin2011; Penarrubia2016; Kravtsov2024). An alternative scenario has the LMC on its second pericentric passage, although most features of the earlier passage are superseded by the more recent passage at a smaller pericenter (Vasiliev2024). Many Local Group phenomena require a large mass for the LMC to explain: for example, the kinematics of its globular clusters (Watkins2024), the kinematics of MW satellites (CorreaMagnus2022; Kravtsov2024); dynamical models of MW stellar streams (2019MNRAS.487.2685E; Koposov2019; Shipp2021; Vasiliev2021; Koposov2023); and the timing argument (Penarrubia2016, but see also Benisty:2022; Chamberlain:2023; Benisty2024) all require an LMC mass $M_{\mathrm{LMC}}\sim 10-20\,\times 10^{10}\,\mathrm{M}_{\odot}$ (see fig. 1 of Vasiliev2023, for a summary of LMC mass estimates). The orbit of the LMC is sensitive to the assumed Galactic potential (see fig. 3 of Vasiliev2023) and, because the LMC is of considerable mass, it is also subject to dynamical friction from the MW dark matter halo (Chandrasekhar1943).
@@
At present-day the LMC is at a Galactocentric distance $d=49.6\pm 0.5\,\mathrm{kpc}$ (Pietrzynski2019), and heliocentric line of sight velocity of $v_{\mathrm{los},}=262.2\pm 3.4\,\mathrm{km}\,\mathrm{s}^{-1}$ (vanderMarel2014). There is more ambiguity in the measurement of the position of the LMC centre (see sec. 4.1, vanderMarel2014) which correspond to different reported values of the mean proper motions (GaiaCollaboration2018; GaiaCollaboration2021; Wan2020) due to both perspective effects and internal motion in the LMC. Table. 2 in Vasiliev2023 gives an overview of some of the recent measurements of the mean LMC proper motions and the associated centre. Although, more recently, hyper velocity stars that are ejected via interactions with the LMC’s supermassive black hole have been used to constrain both the LMC’s dynamical centre and past orbit (Han2025; Lucchini2025).
@@
As the LMC, a satellite galaxy, falls into the gravitational potential of the MW, the central galaxy, responds by generating a density wake (Chandrasekhar1943). This is because an infalling satellite will have a broad range of orbital frequencies that resonate with dark matter particles of the host galaxy (Mulder1983; Weinberg1986). The classical ‘conic’ wake trailing the LMC is described as the transient response, whereas the response elsewhere in the MW halo is the collective response caused by the barycentre displacement (Garavito-Camargo2019; 2021ApJ...919..109G; Tamfal2021; Foote2023). The density wake of the LMC is predicted to leave an observable signature in the density and kinematics of MW halo stars (e.g., Conroy2021; Cavieres2024; Chandra2025; Yaaqib2024; Bystrom2025; Amarante2024).
@@
As the LMC is massive, $M_{\mathrm{LMC}}\sim 10-20\,\times 10^{10}\,\mathrm{M}_{\odot}$ (a 1 : 5–10 mass ratio with the MW), and it has just passed its pericentre with a relative velocity $300\,\mathrm{km}\,\mathrm{s}^{-1}$ (Hunt2025), this has caused significant dynamical disequilibrium in the Galaxy. Specifically, the inner and outer parts of the MW halo have experienced different strengths of gravitational acceleration due to the LMC. To the Galactocentric observer, this appears as if the Galactic Northern sky is red-shifted and the Southern sky is blue-shifted as the halo moves preferentially ‘up’, towards the Galactic North. This displacement, known as the ‘reflex motion’, manifests itself as a dipole signal in density (Belokurov2019; Garavito-Camargo2021; Conroy2021; Amarante2024) that is higher in the Galactic North, and similarly there exists a dipole in stellar radial velocities (2019MNRAS.487.2685E; 2020MNRAS.494L..11P; Petersen2021; Erkal2021; Yaaqib2024; Chandra2025b; Chandra2025; Bystrom2025; Sheng2025). The magnitude of the velocity dipole is called the travel velocity, $v_{\rm{travel}}$, and its orientation is called the apex direction of the reflex motion $(l_{\rm{apex}}$, $b_{\rm{apex}})$ in Galactic coordinates. Although recent studies agree that the travel velocity direction and magnitude should point towards a point on the past orbit of the LMC, they do not converge on their values (e.g., Bystrom2025, fig. 9). This is likely due to each study using different stellar tracers over varying radial ranges and sky coverages.
@@
The MW–LMC system can be modelled at varying levels of simulation fidelity. There exists a trade off between the simulation fidelity and the ability to explore large model parameter spaces (e.g., see, Vasiliev2021, in the context of modelling the Sagittarius stream). The simplest prescription are rigid models of the MW and LMC galaxies. Rigid models describe the MW and LMC as analytic potentials that have fixed functional forms i.e., they are time-invariant, although they are allowed to move in response to each other. The main advantage of these models is their computational inexpense ($\sim\mathcal{O}(\rm seconds)$). However, the main drawback is their inability to truly capture the physics of galaxy formation e.g., the galaxy masses will change over time, which changes the gravitational potential, and hence their resulting interaction. To accommodate many further complexities, the MW–LMC system can instead be modelled using $N$-body simulations in combination with Basis Function Expansions (BFEs, e.g., Lilley2018a; Lilley2018b; Sanders2020; 2020MNRAS.494L..11P; Garavito-Camargo2019; Garavito-Camargo2021; Lilleengen2023; Vasiliev2024). These models are often labelled as idealised or deforming MW–LMC simulations. They aim to match the present-day conditions of the MW and LMC while also accounting for the deformations to both galaxies due to the re-distribution of dark matter due to the LMC’s infall. This includes the formation of a dark matter dynamical friction wake trailing the LMC (Chandrasekhar1943). These deformations are not something that the simpler rigid MW–LMC models can capture although the dynamical friction can still be accounted for. However, they are significantly more computationally expensive to run ($\sim\mathcal{O}(\rm months)$) as they contain billions of particles. As such, the parameter space explored for MW and LMC properties is much smaller e.g., only a handful of models exist to vary the LMC’s mass (Garavito-Camargo2019). Further, there have been MW–LMC pairs identified within hydrodynamical cosmological simulations. For example, there are many cases within the Feedback In Realistic Environments (FIRE, Wetzel2023) Latte, the Milky Way-est (Buch2024), the DREAMS (Rose2025) and the Auriga simulation suites (Grand2017; Grand2024). This is by far the most physically realistic setting to study MW–LMC interactions. However, in a similar vein to deforming simulations, these cosmological simulations can be computationally expensive and cover somewhat limited model parameter spaces.
@@
Simulation Based Inference (SBI, see, Cranmer2020, for a recent review) offers a route to strike a balance between simulation fidelity and the exploration of model parameter space. SBI is a powerful statistical framework for performing inference in complex modelling scenarios where traditional analytic methods are impractical or impossible. It is particularly well-suited for estimating parameter posterior distributions when the likelihood function cannot be explicitly defined, or is intractable from simulation. Instead, forward simulations are used to generate samples of the data and parameters. This technique is also known as ‘likelihood free inference’ or ‘distribution free inference’. Although, these names are slightly misleading as the inference is not likelihood-free, rather we avoid explicitly defining a likelihood and instead model it using many forward simulations. For our application to the dynamics of the MW–LMC interaction, removing the requirement to provide a well-defined likelihood is beneficial as we are no longer biased by assuming some functional form e.g., a Gaussian distribution for the MW reflex parameters: $v_{\rm{travel}}$, $l_{\rm{apex}}$, and $b_{\rm{apex}}$ (Petersen2021; Yaaqib2024; Bystrom2025; Chandra2025). This is especially important in the context of cosmological MW–LMC analogues as both the model parameter space and the posteriors can be non-trivial. SBI has been applied across many fields of research: from mathematics (e.g., Verdier2022), to quantitative finance (e.g., Yu2012), to seismology (e.g., Saoulis2024), therefore owing itself to become an effective lingua franca for facilitating communication across scientific domains. Furthermore, SBI has been increasingly used throughout astrophysics problems (e.g., Weyant2013; Alsing2019; Jeffrey2021; Lemos2021; Hermans2021; vonWietersheim-Kramsta2024; Lovell2024; Widmark2025; Sante2025; XiangyuanMa2025; Jeffrey2025).
@@
For the first time, we will demonstrate the validity of using SBI to infer properties of the MW–LMC system when using measured dynamical quantities, i.e., the reflex motion, from many forward models of the MW–LMC system using rigid prescriptions. We will investigate the ability of our SBI framework to infer the true parameter values of MW–LMC simulations described by rigid, deforming and cosmological models. This will demonstrate the ability that SBI has to upscale the fidelity of simpler simulations to more complex ones. Ultimately, this allows more a rapid exploration of large model parameters spaces at the fraction of the computational cost required to run the higher fidelity simulations. We will show that our SBI framework is reliable and unbiased through a series of diagnostic tests. Our novel approach will be applied to real data with the aim of constraining properties of the MW, LMC and the dynamics of the induced MW reflex motion.
@@ 2 Simulations
This section provides the details for our low fidelity rigid MW–LMC simulations. Throughout this work, conversions between the heliocentric and Galactocentric cartesian frames use the right-handed Cartesian coordinate system with the Sun positioned at $\vec{r}_{\odot}=(-8.3,0.0,0.02)\,\rm{kpc}$ (GRAVITYCollaboration2019; Bennett2019), with velocity $\vec{v}_{\odot}=(11.1,244.24,7.24)\,\rm{km}\,\rm s^{-1}$ (Schonrich2010; Eilers2019).
@@ 2.1 The Milky Way – LMC potentials
We use the galaxy dynamics C++/Python package agama (2019MNRAS.482.1525V) to generate many rigid MW–LMC simulations, each with a unique combination of model parameters. We describe the models for the MW and LMC in the following sections.
@@ 2.1.1 The Milky Way
Our MW model consists of a spherical bulge, an exponential stellar disc, and a dark matter halo. We keep the parameters of the stellar distributions fixed and close to the commonly used values suggested by McMillan2017. The bulge density profile is:
@@
with a total mass $5\times 10^{10}\,\mathrm{M}_{\odot}$, a scale radius $R_{d}=3\,\mathrm{kpc}$, and a scale height $h_{d}=0.4\,\mathrm{kpc}$. We use a Navarro-Frenk-White (NFW, Navarro1996; Navarro1997) dark matter halo density profile described by $M_{200,\rm MW}$ and $c_{200}$. These quantities are defined by a sphere enclosing an overdensity that is 200 times the critical density of the Universe, $\rho_{\mathrm{crit}}=3H^{2}/8\pi G$, as denoted by the ‘$200$’ subscript. We constrain the normalisation of the halo mass profile such that the circular velocity at the solar position is approximately $235\,\mathrm{km}\,\mathrm{s}^{-1}$ (e.g., matching constraints from McMillan2017, within the associated uncertainty). In combination with a chosen $M_{200,\rm MW}$ value, this allows us to find the corresponding concentration parameter, $c_{200}$.
@@ 2.1.2 The LMC
We model the LMC as a Hernquist dark matter halo (Hernquist1990). We impose the observational constraint from vanderMarel2014 that the derived rotation curve peaks at $91.7\pm\,18.8\,\mathrm{km}\,\mathrm{s}^{-1}$ at $8.7\,\mathrm{kpc}$, which corresponds to an enclosed dynamical mass of $M(<8.7\,\mathrm{kpc})=1.7\pm 0.7\times 10^{10}\,\mathrm{M_{\odot}}$. The mass enclosed of a Hernquist profile within a given radius is $\mathrm{M}(<r)=\mathrm{M_{tot}}\,r^{2}/(r+a)^{2}$ (Hernquist1990). Hence, for a given total and enclosed LMC mass, we can solve for, $a$, the scale radius.
@@ 2.1.3 The MW–LMC interaction
where $\mathbf{x}_{i}$, $\mathbf{v}_{i}$ and $\Phi_{i}$ (where $i=\{\mathrm{MW},\mathrm{LMC}\}$) are the position vectors, velocity vectors and the gravitational potentials of the MW and LMC in an inertial frame, respectively. The extra acceleration term, $\mathbf{a}_{\mathrm{DF}}$, accounts for Chandrasekhar dynamical friction on the trajectory of the LMC (Chandrasekhar1943; Binney2008; Jethwa2016),
@@
where $X=v_{\rm LMC}/\sqrt{2}\sigma_{\rm MW}$ and $\sigma_{\rm MW}$ and $\rho_{\rm MW}$ are the velocity dispersion and total density field of the MW. Following Vasiliev2021, we take a fixed value of $\sigma_{\rm MW}=120\,\rm km\,\rm s^{-1}$ for the velocity dispersion as the dynamical friction is insensitive to the precise value. For the Coulomb logarithm we adopt $\ln\Lambda=\ln\left(100\,\rm kpc/\epsilon\right)$. The softening length, $\epsilon$, depends on the satellite’s density profile (White1976). We adopt $\epsilon=1.6\,a_{\rm LMC}$ as this has been used previously when modelling the LMC as a Plummer sphere (e.g., Hashimoto2003; Besla2007; vanderMarel2012; Sohn2013; Kallivayalil2013). The numerator in the Coulomb logarithm expression an arbitrarily chosen value that loosely describes the average separation of the MW and LMC. In principle this value could be updated through the integration of the LMC orbit, however this will have a small effect. Furthermore, our use of a dimensionless parameter, $\lambda_{\mathrm{DF}}$, that modulates the strength of the dynamical friction that the LMC experiences will take into account changes to the fixed Coulomb logarithm value per simulation. Finally, we include the acceleration due to the reflex motion of the MW’s centre of mass towards the LMC.